Use the formula to find the variance and the standard deviation for the given sample data. Round your answer to one more decimal place than the original data. Show all the work.
7 9 12 5 8
Answer #1

Given data - X :- 7,9,12,5,8

Mean = sum / n = 41 / 5 = 8.2

(Xi - mean )² = 1.44, 0.64, 14.44, 10.24, 0.04

SSdif = sum(Xi-mean)² = 26.8

Sample variance = SSdif / n-1 = 26.8/4 = 6.7

Sample standard deviation = sqrt ( variance ) = sqrt(6.7) = 2.588 ~ 2.6